OpenStudy (anonymous):

emily is constructing a triangle. one side is 3 inches long and another is5 inches long. Which of the following is the smallest measure emily could use for the third side of the triangle

OpenStudy (anonymous):

Possible answers?

OpenStudy (anonymous):

Well, if the third side is not the biggest, then \[ 5<3+b \]

OpenStudy (anonymous):

That is the longest side is less than the sum of the other two sides.

OpenStudy (anonymous):

my choices are: 2 inches, 3 inches, 4 inches, or 5 inches

OpenStudy (anonymous):

Okay, in the inequality \[ 5<3+b \]Solve for \(b\).

OpenStudy (wolf1728):

The Triangle Inequality Theorem states that: Any side of a triangle must be: less than the sum of the other 2 sides and greater than the difference of the other 2 sides. sum = 3 + 5 = 8 difference = 5 - 3 = 2 So > 2 third side < 8